OpenStudy (anonymous):

The perimeter of a rectangular sheet is 84cm.If the length of the sheet is increased by 3 cms and breathe is decreased by 1 cm the area of the sheet is increased by 15 cm^2.what is the difference in length and breathe of original rectangular sheet of paper?

OpenStudy (anonymous):

2x+2y=84 x+y=42 x=42-y (x+3)(y-1)=xy+15 xy+3y-x-3=xy+15 3y-x=18 3y-(42-y)=18 4y=60 y=15 x=27

OpenStudy (anonymous):

awesome...!

myininaya (myininaya):

hey kuma look at your previous queston